OpenStudy (anonymous):

simplified form of 15qr5/3qr4

myininaya (myininaya):

\[\frac{15 \sqrt{5}}{3 \sqrt{4}} ?\]

OpenStudy (anonymous):

cubed root maybe?

OpenStudy (anonymous):

in any case \[\frac{15}{3}=5\]

myininaya (myininaya):

well \[\frac{15}{3}=5 \text{ and } \sqrt{4}=2\]

OpenStudy (anonymous):

but if it is cubed root all bets are off

OpenStudy (anonymous):

\[\frac{15\sqrt[3]{5}}{3\sqrt[3]{4}}\] \[\frac{5\sqrt[3]{5}}{\sqrt[3]{4}}\] \[\frac{5\sqrt[3]{5}}{\sqrt[3]{4}}\times\frac{ \sqrt[3]{2}}{\sqrt[3]{2}}\] \[\frac{5\sqrt[3]{10}}{\sqrt[3]{8}}\] \[\frac{5\sqrt[3]{10}}{2}\]

OpenStudy (anonymous):

5r